# Université de Genève, CH-1211 Genève, Switzerland

### v1

@@ 5.1 Evaluation at rational points
[…]
$y_{\pm,k}(x)=y_{\pm}(x)+2\pi{\rm i}\frac{k}{Q}\,,\quad 0\leq k\leq PQ-1\,.$ (5.183)
@@
$\frac{1}{G\left(x,y_{\pm,k}^{(\epsilon)}(x)\right)}=\mp\frac{1}{\epsilon Q\sqrt{\Delta({\rm e}^{xQ})}}+{\cal O}(\epsilon^{0})\,,\quad 0\leq k\leq PQ-1\,.$ (5.184)
Note that the leading-order term in the expansion above does not depend on the integer $k$. The integral in Eq. (5.178) can then be evaluated via the sum of residues of the integrand in Eq. (5.179) at the singularities in Eq. (5.181). We find in this way that
[…]
$=-\frac{2\pi{\rm i}{\rm e}^{2xQ}}{Q\sqrt{\Delta({\rm e}^{xQ})}}\sum_{k=0}^{PQ-1}{\rm e}^{-\frac{2xk}{P}}\left({\rm e}^{F_{k}(y_{+}(x))}-{\rm e}^{F_{k}(y_{-}(x))}\right)\,,$
@@
$F_{k}(y(x))=\frac{{\rm i}Qxy(x)}{\pi P}+\log\Phi_{{\mathsf{b}}}\left(\frac{x-y(x)}{2\pi{\mathsf{b}}}-{\rm i}{\mathsf{b}}\frac{k}{P}\right)-\log\Phi_{{\mathsf{b}}}\left(\frac{y(x)}{2\pi{\mathsf{b}}}+{\rm i}{\mathsf{b}}\frac{k}{P}\right)\,,$ (5.186)
for $10\leq k\leq PQ-$ 1. Therefore, substituting Eq. (5.185) into Eq. (5.170), we obtain that the state-integral invariant of the figure-eight knot at rational values is given exactly by
$\chi_{\bm{4}_{1}}(u;\hbar)=\frac{{\rm e}^{\frac{x^{2}Q}{2\pi{\rm i}P}+x\left(2Q-\frac{1}{2}-\frac{Q}{2P}\right)}}{{\rm i}{\mathsf{b}}Q\sqrt{\Delta({\rm e}^{xQ})}}\sum_{k=0}^{PQ-1}{\rm e}^{-\frac{2xk}{P}}\left({\rm e}^{F_{k}(y_{+}(x))}-{\rm e}^{F_{k}(y_{-}(x))}\right)\,.$ (5.187)
We conclude by noting that $\Phi_{{\mathsf{b}}}(z)$ for ${\mathsf{b}}^{2}\in{\mathbb{Q}}_{>0}$ can be expressed explicitly in terms of more elementary functions garou-kas . The relevant formula is written in Eq. (C.11) and can be used to simplify the computational implementation of Eq. (5.187) for arbitrary $P,Q\in{\mathbb{Z}}_{>0}$ coprime.
Figure 2: The state integral $\chi_{\bm{4}_{1}}(u;\hbar)$ in the closed form of Eq. (5.187) as a function of $x\in{\mathbb{R}}$, at the bottom, and its components $\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar)$ in Eq. (5.188), in the top left and top right, respectively, for $P=Q=1$. The real and imaginary parts are displayed in blue and red.
We will now comment on the most notable features of the exact expression for the non-perturbative wave function in Eq. (5.187). It will be useful to introduce
$\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar)=\frac{{\rm e}^{\frac{x^{2}Q}{2\pi{\rm i}P}+x\left(2Q-\frac{1}{2}-\frac{Q}{2P}\right)}}{{\rm i}{\mathsf{b}}Q\sqrt{\Delta({\rm e}^{xQ})}}\sum_{k=0}^{PQ-1}{\rm e}^{-\frac{2xk}{P}}{\rm e}^{F_{k}(y_{\pm}(x))}\,,$ (5.188)
@@
Observe that the $S$-duality, which acts in the rational case via $P\leftrightarrow Q$ and $x\rightarrow xQ/P$, leaves $\Delta({\rm e}^{xQ})$ unchanged while transforming $y_{\pm,k}(x)$ into $y_{\pm,k}(x)Q/P$, and therefore the closed formula for the state integral of the $\bm{4}_{1}$-knot in Eq. (5.187) is $S$-invariant, as expected. Similarly, we also find that its components in Eq. (5.188) are themselves invariant under $S$-duality. Moreover, we have that
$\Im(\chi_{\bm{4}_{1}}^{(+)}(u;\hbar))=\Im(\chi_{\bm{4}_{1}}^{(-)}(u;\hbar))\,,$ (5.190)
@@
while both $\Re(\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar))$ and $\Im(\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar))$ are even functions of $x\in{\mathbb{R}}$ with singularities located at the points
$x=\frac{1}{Q}\log\left(\frac{3\pm\sqrt{5}}{2}\right)\,,$ (5.191)
@@
which correspond to the zeros of $\Delta({\rm e}^{xQ})$ in Eq. (5.182) and, as such, do not depend on the value of $P$. The discontinuities of $\Re(\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar))$ cancel with each other to give a smooth difference function in Eq. (5.189), which is itself symmetric under $x\rightarrow-x$. Remarkably, we find that the total state integral $\chi_{\bm{4}_{1}}(u;\hbar)$ is a real-valued smooth square-integrable function of $x\in{\mathbb{R}}$, while both functions $\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar)$ are neither real-valued, smooth, or square-integrable when considered separately. To illustrate these properties with an example, we show in Fig. 2 the functions in Eqs. (5.187), (5.188) for the benchmark choice $P=Q=1$. The same features and similar functional profiles are verified for various choices of $P,Q\in{\mathbb{Z}}_{>0}$ coprime. We show in Fig. 3 the full non-perturbative wave function for a sample of small values of ${\mathsf{b}}^{2}\in{\mathbb{Q}}_{>0}$.
@@
Figure 3: The real part of the state integral $\chi_{\bm{4}_{1}}(u;\hbar)$ in the closed form of Eq. (5.187) as a function of $x\in{\mathbb{R}}$ for rational values of ${\mathsf{b}}^{2}$. We show ${\mathsf{b}}^{2}=1/3$ (in red), $1/2$ (in green), $2/3$ (in blue), and $1$ (in black), on the left, and ${\mathsf{b}}^{2}=3$ (in red), $2$ (in green), $3/2$ (in blue), and $1$ (in black), on the right. The imaginary part is identically zero and not displayed explicitly in the plot.
@@
We observe that $\chi_{\bm{4}_{1}}(u;\hbar)$ has a Gaussian-type profile centered at $x=0$ with lateral oscillations around the $x$-axis, which are quickly dampened. The peak at the origin can be evaluated directly from Eqs. (5.181), (5.186), and (5.187) for arbitrary $P,Q\in{\mathbb{Z}}_{>0}$ coprime, yielding a result which is left unchanged by $P\leftrightarrow Q$, as expected from the $S$-invariance of the state integral and evident in Fig. 3. In particular, the highest peak corresponds to the choice $P=Q=1$ and amounts to
$\chi_{\bm{4}_{1}}(0;2\pi)=\frac{2}{\sqrt{3}}\sinh\left(\frac{V}{2\pi}\right)\,,\quad V=2\,\Im\left(\text{Li}_{2}\left({\rm e}^{\frac{\pi{\rm i}}{3}}\right)\right)=2.0298832\dots\,,$ (5.192)
@@
where $\text{Li}_{2}(z)$ is the standard dilogarithm and $V$ is the volume of the $\bm{4}_{1}$-knot, as expected garou-kas . As ${\mathsf{b}}^{2}>1$ increases, the shape of the function $\chi_{\bm{4}_{1}}(u;\hbar)$ widens, while for decreasing ${\mathsf{b}}^{2}<1$, the shape tightens, and in both cases the peak lowers.
@@
We point out that the structure of the exact wave function at rational values for the figure-eight knot obtained here is tantalizingly close to what has been found in the spectral theory of quantum mirror curves. The case of $\hbar=2\pi$, corresponding to ${\mathsf{b}}^{2}=1$, was studied in detail in mz-wv ; mz-wv2 . As in Eq. (5.189), the wave functions in mz-wv ; mz-wv2 are the sums of two individual functions corresponding to two different choices of branch cut in the mirror curve. Each component function is singular when considered alone. Still, the singularities cancel each other in the sum, and one obtains a square-integrable wave function in the end, similar to the one plotted in Fig. 2. We observe that, in the case of the quantum mirror curves studied in mz-wv ; mz-wv2 , one is working with exponentiated Heisenberg operators, and the resulting wave functions do not satisfy the nodal theorem for the Schrödinger operator typical of confining potentials, since the ground state changes sign in a small region. This same feature is also evident in our plots for the exact wave functions of the $\bm{4}_{1}$-knot at rational points shown in Fig. 2 and Fig. 3. Such a property suggests that the wave function occurring in complex CS theory is a ground state for an appropriate spectral definition of the quantum $A$-polynomial. We will return to this point in the conclusions.
@@
Finally, let us stress that the results above provide a third method for the computation of the exact wave function at rational values of $\hbar$ to be compared with the previous two presented in Section 3. We conclude by verifying numerically that our closed formula in Eq. (5.187) for the state integral of the figure-eight knot at rational values matches the results of Section 4.1. Specifically, the two components $\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar)$ in Eq. (5.188) can be identified with the two terms in the holomorphic/antiholomorphic block decomposition in Eq. (4.134), which correspond to the conjugate and geometric classical branches, respectively. This identification will be formally clarified in the discussion in the following section.
@@ 5.2 Factorization and two special subcases
Let us apply the explicit formula for the Faddeev’s quantum dilogarithm at the rational point ${\mathsf{b}}^{2}=P/Q$, $P,Q\in{\mathbb{Z}}_{>0}$ coprime, in Eq. (C.11) to the expression for $F_{k}(y(x))$, $10\leq k\leq PQ-$ 1, which is given in Eq. (5.186). After appropriately rearranging the terms, we find the factorization
@@
$f_{0,k}(y(x))$ $=\left[\left(1-{\rm e}^{Q(x-y(x))+\pi{\rm i}(P+Q)}\right)\left(1-{\rm e}^{Qy(x)+\pi{\rm i}(P+Q)}\right)\right]^{\frac{k}{PQ}}\,,$ (5.194a)
$f_{1,k}(y(x))$ $=\prod_{j=1}^{Q-1}\left[\frac{\left(1+{\rm e}^{x-y(x)+\frac{\pi{\rm i}P}{Q}(2j+1)}\right)\left(1+{\rm e}^{y(x)+\frac{\pi{\rm i}P}{Q}(2j+1+2k/P)}\right)}{\left(1+{\rm e}^{x-y(x)+\frac{\pi{\rm i}P}{Q}(2j+1-2k/P)}\right)\left(1+{\rm e}^{y(x)+\frac{\pi{\rm i}P}{Q}(2j+1)}\right)}\right]^{j\over Q}\,,$ (5.194b)
$f_{2,k}(y(x))$ $=\prod_{j=1}^{P-1}\left[\frac{\left(1+{\rm e}^{\frac{Q}{P}(x-y(x))+\frac{\pi{\rm i}Q}{P}(2j+1)}\right)\left(1+{\rm e}^{\frac{Q}{P}y(x)+\frac{\pi{\rm i}Q}{P}(2j+1+2k/Q)}\right)}{\left(1+{\rm e}^{\frac{Q}{P}(x-y(x))+\frac{\pi{\rm i}Q}{P}(2j+1-2k/Q)}\right)\left(1+{\rm e}^{\frac{Q}{P}y(x)+\frac{\pi{\rm i}Q}{P}(2j+1)}\right)}\right]^{j\over P}\,,$ (5.194c)
for $10\leq k\leq PQ-$ 1. Thus, we can write the state integral for arbitrary rational values of ${\mathsf{b}}^{2}$ in Eq. (5.187) in the equivalent form
$\chi_{\bm{4}_{1}}(u;\hbar)=\frac{{\rm e}^{\frac{x^{2}Q}{2\pi{\rm i}P}+x\left(2Q-\frac{1}{2}-\frac{Q}{2P}\right)}}{{\rm i}{\mathsf{b}}Q\sqrt{\Delta({\rm e}^{xQ})}}\left[{\rm e}^{F_{0}(y_{+}(x))}{\cal A}(y_{+}(x))-{\rm e}^{F_{0}(y_{-}(x))}{\cal A}(y_{-}(x))\right]\,,$ (5.195)
@@
${\cal A}(y(x))=\sum_{k=0}^{PQ-1}{\rm e}^{-\frac{2xk}{P}}f_{0,k}(y(x))f_{1,k}(y(x))f_{2,k}(y(x))\,,$ (5.196)
and $F_{0}(y(x))$ can be expressed explicitly in terms of more elementary functions using Eq. (C.11). We stress that the formula for the state integral in Eq. (5.195) allows us to more clearly separate the contributions from the two non-abelian saddle connections of the $\bm{4}_{1}$-knot labeled by $\alpha=\text{geom},\text{conj}$, as in the holomorphic/antiholomorphic block decomposition in Eq. (3.51). Indeed, expanding the functions $F_{0}(y_{\mp}(x))$ in powers of $X,X^{-1}$ and suitably identifying the terms in the expansions, we derive the integer constants $c_{\pm s}$, $s\in{\mathbb{Z}}_{>0}$, previously written in Eqs. (4.144a), (4.144b), for the geometric and conjugate branches, respectively.
Let us consider now the two special choices ${\mathsf{b}}^{2}=P$ and ${\mathsf{b}}^{2}=1/Q$ for $P,Q\in{\mathbb{Z}}_{>0}$. In both subcases, the closed formula for the state integral of the figure-eight knot in Eq. (5.195) specializes in a correspondingly simpler and more compact form. We start by choosing ${\mathsf{b}}^{2}=P$ positive integer. After some manipulations, the factorization formula in Eq. (5.193) reduces to the relation
${\rm e}^{F_{k}(y(x))}={\rm e}^{F_{0}(y(x))}\prod_{j=0}^{k-1}\left(1+{\rm e}^{\frac{1}{P}(x-y(x))-\frac{\pi{\rm i}}{P}(2j+1)}\right)\left(1+{\rm e}^{\frac{1}{P}y(x)+\frac{\pi{\rm i}}{P}(2j+1)}\right)\,,$ (5.197)
where $10\leq k\leq P-$ 1, and we can write the state integral in Eq. (5.195) for ${\mathsf{b}}^{2}=P$ as
$\chi_{\bm{4}_{1}}(u;\hbar)=\frac{{\rm e}^{\frac{x^{2}}{2\pi{\rm i}P}+x\left(\frac{3}{2}-\frac{1}{2P}\right)}}{{\rm i}{\mathsf{b}}\sqrt{\Delta({\rm e}^{x})}}\left[{\rm e}^{F_{0}(y_{+}(x))}{\cal B}(y_{+}(x))-{\rm e}^{F_{0}(y_{-}(x))}{\cal B}(y_{-}(x))\right]\,,$ (5.198)
@@
${\cal B}(y(x))=\sum_{k=0}^{P-1}{\rm e}^{-\frac{2xk}{P}}\prod_{j=0}^{k-1}\left(1+{\rm e}^{\frac{1}{P}(x-y(x))-\frac{\pi{\rm i}}{P}(2j+1)}\right)\left(1+{\rm e}^{\frac{1}{P}y(x)+\frac{\pi{\rm i}}{P}(2j+1)}\right)\,,$ (5.199)
and $F_{0}(y(x))$ can be expressed explicitly by means of Eq. (C.13a). Let us now choose ${\mathsf{b}}^{-2}=Q$ positive integer. After some manipulations, the factorization formula in Eq. (5.193) becomes
${\rm e}^{F_{k}(y(x))}={\rm e}^{F_{0}(y(x))}\prod_{j=0}^{k-1}\left(1+{\rm e}^{x-y(x)-\frac{\pi{\rm i}}{Q}(2j+1)}\right)\left(1+{\rm e}^{y(x)+\frac{\pi{\rm i}}{Q}(2j+1)}\right)\,,$ (5.200)
and the exact partition function in Eq. (5.195) for ${\mathsf{b}}^{2}=1/Q$ assumes the simplified form
$\chi_{\bm{4}_{1}}(u;\hbar)=\frac{{\rm e}^{\frac{x^{2}Q}{2\pi{\rm i}}+x\left(\frac{3Q}{2}-\frac{1}{2}\right)}}{{\rm i}{\mathsf{b}}Q\sqrt{\Delta({\rm e}^{xQ})}}\left[{\rm e}^{F_{0}(y_{+}(x))}{\cal C}(y_{+}(x))-{\rm e}^{F_{0}(y_{-}(x))}{\cal C}(y_{-}(x))\right]\,,$ (5.201)

### v2

@@ 5.1 Evaluation at rational points
and the branch of the logarithm is chosen in such a way that $\operatorname{Im}(y_{\pm,0}(x))\in(0,2\pi/Q]$ for each value of $x\in{\mathbb{R}}$. Note that $\Delta(X)$ is the same function introduced in equation (2.29b), which appears in the formula for the solutions to the classical $A$-polynomial of the figure-eight knot in equation (2.28) and in the formula for the function $\varphi_{Q}(X)$ obtained via the matrix formalism of Section 3.4.2 in equation (4.19). Besides, the functions $y_{\pm,k}(x)$ only depend on $x$ through ${\rm e}^{xQ}=X^{Q}$}. We denote $y_{\pm}(x)=y_{\pm,0}(x)$ for simplicity, so that
$y_{\pm,k}(x)=y_{\pm}(x)+2\pi{\rm i}{{k\over Q}},\qquad 0\leq k\leq PQ-1.$ (5.14)
@@
${{1\over G\bigl(x,y_{\pm,k}^{(\epsilon)}(x)\bigr)}}=\mp{{1\over\epsilon Q\sqrt{\Delta\bigl({\rm e}^{xQ}\bigr)}}}+{\cal O}\bigl(\epsilon^{0}\bigr),\qquad 0\leq k\leq PQ-1.$ (5.15)
Note that the leading-order term in the expansion above does not depend on the integer $k$. The integral in equation (5.9) can then be evaluated via the sum of residues of the integrand in equation (5.10) at the singularities in equation (5.12). We find in this way that
$\int_{{\mathbb{R}}+{\rm i}0}f(y){\rm d}y$ $=2\pi{\rm i}\sum_{k=0}^{PQ-1}\underset{y(x)=y_{\pm,k}(x)}{\operatorname{Res}}\left({{f(y){\rm e}^{2xQ}\over G(x,y)}}\right)$
$=-{{2\pi{\rm i}{\rm e}^{2xQ}\over Q\sqrt{\Delta\bigl({\rm e}^{xQ}\bigr)}}}\sum_{k=0}^{PQ-1}{\rm e}^{-{{2xk\over P}}}\bigl({\rm e}^{F_{k}(y_{+}(x))}-{\rm e}^{F_{k}(y_{-}(x))}\bigr),$ (5.16)
@@
$F_{k}(y(x))={{{\rm i}Qxy(x)\over\pi P}}+\log\Phi_{{\mathsf{b}}}\left({{x-y(x)\over 2\pi{\mathsf{b}}}}-{\rm i}{\mathsf{b}}{{k\over P}}\right)-\log\Phi_{{\mathsf{b}}}\left({{y(x)\over 2\pi{\mathsf{b}}}}+{\rm i}{\mathsf{b}}{{k\over P}}\right),$ (5.17)
for $0\leq k\leq PQ-1$. Therefore, substituting equation (5.16) into equation (5.1), we obtain that the state-integral invariant of the figure-eight knot at rational values is given exactly by
$\chi_{\bm{4}_{1}}(u;\hbar)={{{\rm e}^{{{x^{2}Q\over 2\pi{\rm i}P}}+x\left(2Q-{{1\over 2}}-{{Q\over 2P}}\right)}\over{\rm i}{\mathsf{b}}Q\sqrt{\Delta\bigl({\rm e}^{xQ}\bigr)}}}\sum_{k=0}^{PQ-1}{\rm e}^{-{{2xk\over P}}}\bigl({\rm e}^{F_{k}(y_{+}(x))}-{\rm e}^{F_{k}(y_{-}(x))}\bigr).$ (5.18)
We conclude by noting that $\Phi_{{\mathsf{b}}}(z)$ for ${\mathsf{b}}^{2}\in{\mathbb{Q}}_{>0}$ can be expressed explicitly in terms of more elementary functions [18]. The relevant formula is written in equation (C.18) and can be used to simplify the computational implementation of equation (5.18) for arbitrary $P,Q\in{\mathbb{Z}}_{>0}$ coprime.
Figure 2: The state integral $\chi_{\bm{4}_{1}}(u;\hbar)$ in the closed form of equation (5.18) as a function of $x\in{\mathbb{R}}$, at the bottom, and its components $\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar)$ in equation (5.19), in the top left and top right, respectively, for $P=Q=1$. The real and imaginary parts are displayed in blue and red.
We will now comment on the most notable features of the exact expression for the non-perturbative wave function in equation (5.18). It will be useful to introduce
$\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar)={{{\rm e}^{{{x^{2}Q\over 2\pi{\rm i}P}}+x\left(2Q-{{1\over 2}}-{{Q\over 2P}}\right)}\over{\rm i}{\mathsf{b}}Q\sqrt{\Delta\bigl({\rm e}^{xQ}\bigr)}}}\sum_{k=0}^{PQ-1}{\rm e}^{-{{2xk\over P}}}{\rm e}^{F_{k}(y_{\pm}(x))},$ (5.19)
@@
Observe that the $S$-duality, which acts in the rational case via $P\leftrightarrow Q$ and $x\leftrightarrow xQ/P$, leaves $\Delta\bigl({\rm e}^{xQ}\bigr)$ unchanged while transforming $y_{\pm,k}(x)$ into $y_{\pm,k}(x)Q/P$, and therefore the closed formula for the state integral of the $\bm{4}_{1}$-knot in equation (5.18) is $S$-invariant, as expected. Similarly, we also find that its components in equation (5.19) are themselves invariant under $S$-duality. Moreover, we have that
$\operatorname{Im}\bigl(\chi_{\bm{4}_{1}}^{(+)}(u;\hbar)\bigr)=\operatorname{Im}\bigl(\chi_{\bm{4}_{1}}^{(-)}(u;\hbar)\bigr),$ (5.21)
@@
while both $\operatorname{Re}\bigl(\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar)\bigr)$ and $\operatorname{Im}\bigl(\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar)\bigr)$ are even functions of $x\in{\mathbb{R}}$ with singularities located at the points
$x={{1\over Q}}\log\left({{3\pm\sqrt{5}\over 2}}\right),$ (5.22)
@@
which correspond to the zeros of $\Delta\bigl({\rm e}^{xQ}\bigr)$ in equation (5.13) and, as such, do not depend on the value of $P$. The discontinuities of $\operatorname{Re}\bigl(\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar)\bigr)$ cancel with each other to give a smooth difference function in equation (5.20), which is itself symmetric under $x\rightarrow-x$. Remarkably, we find that the total state integral $\chi_{\bm{4}_{1}}(u;\hbar)$ is a real-valued smooth square-integrable function of $x\in{\mathbb{R}}$}, while both functions $\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar)$ are neither real-valued, smooth, or square-integrable when considered separately. To illustrate these properties with an example, we show in Figure 2 the functions in equations (5.18) and (5.19) for the benchmark choice $P=Q=1$. The same features and similar functional profiles are verified for various choices of $P,Q\in{\mathbb{Z}}_{>0}$ coprime. We show in Figure 3 the full non-perturbative wave function for a sample of small values of ${\mathsf{b}}^{2}\in{\mathbb{Q}}_{>0}$.
@@
Figure 3: The real part of the state integral $\chi_{\bm{4}_{1}}(u;\hbar)$ in the closed form of equation (5.18) as a function of $x\in{\mathbb{R}}$ for rational values of ${\mathsf{b}}^{2}$. We show ${\mathsf{b}}^{2}=1/3$ (in red), $1/2$ (in green), $2/3$ (in blue), and 11 (in black), on the left, and ${\mathsf{b}}^{2}=3$ (in red), 22 (in green), $3/2$ (in blue), and 11 (in black), on the right. The imaginary part is identically zero and not displayed explicitly in the plot.
@@
We observe that $\chi_{\bm{4}_{1}}(u;\hbar)$ has a Gaussian-type profile centered at $x=0$ with lateral oscillations around the $x$-axis that are quickly dampened. The peak at the origin can be evaluated directly from equations (5.12), (5.17), and (5.18) for arbitrary $P,Q\in{\mathbb{Z}}_{>0}$ coprime, yielding a result which is left unchanged by $P\leftrightarrow Q$, as expected from the $S$-invariance of the state integral and evident in Figure 3. In particular, the highest peak corresponds to the choice $P=Q=1$ and amounts to
$\chi_{\bm{4}_{1}}(0;2\pi)={{2\over\sqrt{3}}}\sinh\left({{V\over 2\pi}}\right),\qquad V=2\operatorname{Im}\bigl(\operatorname{Li}_{2}\bigl({\rm e}^{{{\pi{\rm i}\over 3}}}\bigr)\bigr)=2.0298832\dots,$ (5.23)
@@
where $\operatorname{Li}_{2}(z)$ is the standard dilogarithm and $V$ is the volume of the $\bm{4}_{1}$-knot, as expected from [18]. As ${\mathsf{b}}^{2}<1$ decreases, the shape of the function $\chi_{\bm{4}_{1}}(u;\hbar)$ tightens, while for increasing ${\mathsf{b}}^{2}>1$}, the shape widens, and in both cases the peak lowers.
@@
We point out that the structure of the exact wave function at rational values for the figure-eight knot obtained here is tantalizingly close to what has been found in the spectral theory of quantum mirror curves. The case of $\hbar=2\pi$, corresponding to ${\mathsf{b}}^{2}=1$, was studied in detail in [38, 39]. As in equation (5.20), the wave functions in [38, 39] are the sums of two individual functions corresponding to two different choices of branch cut in the mirror curve. Each component function is singular when considered alone. Still, the singularities cancel each other in the sum, and one obtains a square-integrable wave function in the end, similar to the one plotted in Figure 2. We observe that, in the case of the quantum mirror curves studied in [38, 39], one is working with exponentiated Heisenberg operators, and the resulting wave functions do not satisfy the nodal theorem for the Schrödinger operator typical of confining potentials, since the ground state changes sign in a small region. This same feature is also evident in our plots for the exact wave functions of the $\bm{4}_{1}$-knot at rational points shown in Figures 2 and 3. Such a property suggests that the wave function occurring in complex CS theory is a ground state for an appropriate spectral definition of the quantum $A$-polynomial. We will return to this point in the conclusions.
@@
Finally, let us stress that the results above provide a third method for the computation of the exact wave function at rational values of $\hbar$ to be compared with the previous two presented in Section 3. We conclude by verifying numerically that our closed formula in equation (5.18) for the state integral of the figure-eight knot at rational values matches the results of Section 4.1. Specifically, the two components $\chi_{\bm{4}_{1}}^{(\pm)}(u;\hbar)$ in equation (5.19) can be formally identified with the two terms in the holomorphic/antiholomorphic block decomposition in equation (4.1), which correspond to the conjugate and geometric classical branches, respectively. This identification will be clarified in the discussion in the following section.
@@ 5.2 Factorization and two special subcases
Let us apply the explicit formula for the Faddeev’s quantum dilogarithm at the rational point ${\mathsf{b}}^{2}=P/Q$, $P,Q\in{\mathbb{Z}}_{>0}$ coprime, in equation (C.18) to the expression for $F_{k}(y(x))$, $0\leq k\leq PQ-1$, which is given in equation (5.17). After appropriately rearranging the terms, we obtain the factorization
@@
$f_{0,k}(y(x))$ $=\big[\bigl(1-{\rm e}^{Q(x-y(x))+\pi{\rm i}(P+Q)}\bigr)\bigl(1-{\rm e}^{Qy(x)+\pi{\rm i}(P+Q)}\bigr)\big]^{{{k\over PQ}}},$ (5.25a)
$f_{1,k}(y(x))$ $=\prod_{j=1}^{Q-1}\left[{{\bigl(1+{\rm e}^{x-y(x)+{{\pi{\rm i}P\over Q}}(2j+1)}\bigr)\bigl(1+{\rm e}^{y(x)+{{\pi{\rm i}P\over Q}}(2j+1+2k/P)}\bigr)\over\bigl(1+{\rm e}^{x-y(x)+{{\pi{\rm i}P\over Q}}(2j+1-2k/P)}\bigr)\bigl(1+{\rm e}^{y(x)+{{\pi{\rm i}P\over Q}}(2j+1)}\bigr)}}\right]^{j\over Q},$ (5.25b)
$f_{2,k}(y(x))$ $=\prod_{j=1}^{P-1}\left[{{\bigl(1+{\rm e}^{{{Q\over P}}(x-y(x))+{{\pi{\rm i}Q\over P}}(2j+1)}\bigr)\bigl(1+{\rm e}^{{{Q\over P}}y(x)+{{\pi{\rm i}Q\over P}}(2j+1+2k/Q)}\bigr)\over\bigl(1+{\rm e}^{{{Q\over P}}(x-y(x))+{{\pi{\rm i}Q\over P}}(2j+1-2k/Q)}\bigr)\bigl(1+{\rm e}^{{{Q\over P}}y(x)+{{\pi{\rm i}Q\over P}}(2j+1)}\bigr)}}\right]^{j\over P},$ (5.25c)
for $0\leq k\leq PQ-1$. Thus, we can write the state integral for arbitrary rational values of ${\mathsf{b}}^{2}$ in equation (5.18) in the equivalent form
$\chi_{\bm{4}_{1}}(u;\hbar)={{{\rm e}^{{{x^{2}Q\over 2\pi{\rm i}P}}+x(2Q-{{1\over 2}}-{{Q\over 2P}})}\over{\rm i}{\mathsf{b}}Q\sqrt{\Delta\bigl({\rm e}^{xQ}\bigr)}}}\big[{\rm e}^{F_{0}(y_{+}(x))}{\cal A}(y_{+}(x))-{\rm e}^{F_{0}(y_{-}(x))}{\cal A}(y_{-}(x))\big],$ (5.26)
@@
${\cal A}(y(x))=\sum_{k=0}^{PQ-1}{\rm e}^{-{{2xk\over P}}}f_{0,k}(y(x))f_{1,k}(y(x))f_{2,k}(y(x)),$ (5.27)
and $F_{0}(y(x))$ can be expressed explicitly in terms of more elementary functions using equation (C.18). We stress that the formula for the state integral in equation (5.26) allows us to more clearly separate the contributions from the two non-abelian flat $\mathrm{SL}(2,{\mathbb{C}})$-connections of the $\bm{4}_{1}$-knot labeled by $\alpha={\rm geom},\mathrm{conj}$, as in the holomorphic/antiholomorphic block decomposition in equation (3.13). Indeed, expanding the functions $F_{0}(y_{\mp}(x))$ in powers of $X,X^{-1}$ and suitably identifying the terms in the expansions, we derive the integer constants $c^{(\alpha)}_{\pm s}$, $s\in{\mathbb{Z}}_{>0}$, previously written in equations (4.11a) and (4.11b), for the geometric and conjugate branches, respectively.
Let us consider now the two special choices ${\mathsf{b}}^{2}=P$ and ${\mathsf{b}}^{2}=1/Q$ for $P,Q\in{\mathbb{Z}}_{>0}$. In both subcases, the closed formula for the state integral of the figure-eight knot in equation (5.26) specializes in a correspondingly simpler and more compact form. We start by choosing ${\mathsf{b}}^{2}=P$ positive integer. After some manipulations, the factorization formula in equation (5.24) reduces to the relation
${\rm e}^{F_{k}(y(x))}={\rm e}^{F_{0}(y(x))}\prod_{j=0}^{k-1}\bigl(1+{\rm e}^{{{1\over P}}(x-y(x))-{{\pi{\rm i}\over P}}(2j+1)}\bigr)\bigl(1+{\rm e}^{{{1\over P}}y(x)+{{\pi{\rm i}\over P}}(2j+1)}\bigr),$ (5.28)
where $0\leq k\leq P-1$, and we can write the state integral in equation (5.26) for ${\mathsf{b}}^{2}=P$ as
$\chi_{\bm{4}_{1}}(u;\hbar)={{{\rm e}^{{{x^{2}\over 2\pi{\rm i}P}}+x({{3\over 2}}-{{1\over 2P}})}\over{\rm i}{\mathsf{b}}\sqrt{\Delta({\rm e}^{x})}}}\big[{\rm e}^{F_{0}(y_{+}(x))}{\cal B}(y_{+}(x))-{\rm e}^{F_{0}(y_{-}(x))}{\cal B}(y_{-}(x))\big],$ (5.29)
@@
${\cal B}(y(x))=\sum_{k=0}^{P-1}{\rm e}^{-{{2xk\over P}}}\prod_{j=0}^{k-1}\bigl(1+{\rm e}^{{{1\over P}}(x-y(x))-{{\pi{\rm i}\over P}}(2j+1)}\bigr)\bigl(1+{\rm e}^{{{1\over P}}y(x)+{{\pi{\rm i}\over P}}(2j+1)}\bigr),$ (5.30)
and $F_{0}(y(x))$ can be expressed explicitly by means of equation (C.20a). Let us now choose ${\mathsf{b}}^{-2}=Q$ positive integer. After some manipulations, the factorization formula in equation (5.24) becomes
${\rm e}^{F_{k}(y(x))}={\rm e}^{F_{0}(y(x))}\prod_{j=0}^{k-1}\bigl(1+{\rm e}^{x-y(x)-{{\pi{\rm i}\over Q}}(2j+1)}\bigr)\bigl(1+{\rm e}^{y(x)+{{\pi{\rm i}\over Q}}(2j+1)}\bigr),$ (5.31)
and the exact partition function in equation (5.26) for ${\mathsf{b}}^{2}=1/Q$ assumes the simplified form
$\chi_{\bm{4}_{1}}(u;\hbar)={{{\rm e}^{{{x^{2}Q\over 2\pi{\rm i}}}+x({{3Q\over 2}}-{{1\over 2}})}\over{\rm i}{\mathsf{b}}Q\sqrt{\Delta\bigl({\rm e}^{xQ}\bigr)}}}\big[{\rm e}^{F_{0}(y_{+}(x))}{\cal C}(y_{+}(x))-{\rm e}^{F_{0}(y_{-}(x))}{\cal C}(y_{-}(x))\big],$ (5.32)